OpenStudy (anonymous):

Evaluate P(-1/2) if P(x) = x^4 + 3x^2 + 2

OpenStudy (anonymous):

\[p(-1/2)=(-1/2)^{4}+3(-1/2)^{2}+2\] \[p(-1/2)=19/16\]

OpenStudy (anonymous):

sorry \[p(-1/2)=45/16\]

OpenStudy (anonymous):

thank you